Operation Summary for June 11, 2010 - June 14, 2010
Friday June 11
The day shift began with Operations monitoring Tevatron (TeV) store 7886 (320.49E30) and the Recycler stash (130.2E10), with the Antiproton Source (Pbar) stacking (18.5E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.
During the day shift, Recycler personnel accessed MI-31 to reset the Pelletron blower. Operations established beam to the Meson FTTB T-977 experiment. Pbar personnel accessed the upper vault area to make adjustments for the JASMIN (T-994) experiment. Operations reported that a connecting rod bearing for the TeV sector E2 wet engine had failed; cryo system technicians made temporary repairs. MI personnel conducted MI-8 collimator studies and caused vacuum pumps to trip off.
During the evening shift, Operations disabled the T-977 beam. Operations terminated store 7886 at 7:20 PM, put the TeV into shot setup, and established store 7887 at 8:45 PM with an average initial luminosity (AIL) of 344E30. Operators monitored the E2 wet engine.
Saturday June 12
The midnight shift began with Operations monitoring store 7887 (184.68E30) and the stash (111.2E10), with Pbar stacking (17.2E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations established beam to T-977 at 4:15 AM. Operations reported several Linac trips due to power amplifier crowbars and a gun spark.
During the day shift, Operations terminated store 7887 at 8:34 AM. Cryo system techs began wet engine repairs at 9:04 AM that lasted the whole shift. Operators prepared CDF, TeV A-E, and F sector for accesses to repair vacuum pumps. Feeder 97 tripped off on an “Instantaneous Neutral” indication; FESS personnel investigated and then reset the breaker.
During the evening shift, Operation began the TeV turn on and cooldown. Operations disabled the T-977 beam at 6:15 PM. Operations reported short downtimes due to two Linac RF stations (LRF2 & 4) and to a MI RF station (MIRF6). Operations put the TeV into shot setup at 9:46 PM and established store 7889 at 10:43 PM with an AIL of 336.9E30.
Sunday June 13
The midnight shift began with Operations monitoring store 7889 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
During the midnight shift, The H- Source output began to drop; operators raised the source pressure. Two P1 transfer line ion pumps tripped off; they reset. Operations reported short downtimes due to minor RF trips for different machines.
During the day shift, Operations terminated store 7889 at 9:30 AM, put the TeV into shot setup, and established store 7890 at 10:58 AM with an AIL of 346.3E30. MIRF1 tripped off during shot setup and would not reset; operators bypassed the station.
During the evening shift, a network switch card failed at 6:30 PM that caused the loss of communication with various systems on site; Controls personnel had this fixed by 8 PM.
Monday June 14
The midnight shift began with Operations monitoring store 7890 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 7890 at 1:29 AM, put the TeV into shot setup, and established store 7891 at 2:43 AM with an AIL of 312E30. Operations established beam to T-977 around 4 AM.
The Plans for Monday
The plans for today are to stack, stash, and store.
Comments and Suggestions
|last modified 06/14/2010 email Fermilab|
|Security, Privacy, Legal|